# Pushing the limits of eccentricity in planet-disc interactions

### v1

@@ 5.2.2 Higher mass planets
Increasing the planetary mass further would result in more pronounced differences which we do not consider in this paper. Zhu & Zhang (2022) performed exemplar 2D simulations of eccentric Jupiter mass planets, showing that the spiral wake morphology is significantly modified compared with the low mass regime. Namely, the larger amplitude wakes can interact nonlinearly, leading to the messier interference of wake crossings. Furthermore, the rapid steepening and shocking of the waves deposits angular momentum more efficiently. This can promote gap opening and a transition towards Type II migration. Interestingly, as the eccentricity increases, the gap depth is found to be shallower (also see Duffell & Chiang, 2015) which might help constrain the properties of hypothesized planets in observed gaps.
@@ 5.2.4 Higher eccentricities
Whilst we push to highly supersonic values of $e$, we curtail our study at the upper value of $e=0.7$. By this point we are beginning to find small but noticeable discrepancies in the linear theory and simulation results. For example, the amplitude of sharp radial torque density features are not in exact agreement in Figs. 3-4 for $e=0.5$, whilst the linear theory seems to slightly underestimate the magnitude of numerical net torques in Fig. 6. However, in appendix A.1 we have demonstrated sufficient modal convergence in $|m-l|$ for the fiducial upper value of $m=170$ adopted by the linear framework. We also tested the use of expanded ranges with $m_{\rm max}=220$ and $|m-l|_{\rm max}=80$, but this did not provide a significant improvement on the comparison with simulations.
@@ Acknowledgements
CWF and AJD would like to acknowledge insightful conversations with Roman Rafikov, Yoram Lithwick, and Eugene Chiang during the preparation of this manuscript. We also thank Pablo Benítez-Llambay for suggesting the inclusion of appendix B as a benchmark for future numerical investigations. This research was supported by the W. M. Keck Foundation Fund and the IAS Fund for Memberships in the Natural Sciences (CWF), and by NASA through the NASA Hubble Fellowship grant #HST-HF2-51553.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555 (AJD).
@@ A.1 Linear theory modal convergence
In this section, we will interrogate the modal convergence of our linear methodology. Indeed, since this paper considers highly eccentric orbits, beyond the scope which was confidently converged in the previous work of FR25, we anticipate the need for a high number of modes. Throughout the main body of this paper, we have assumed a default modal range of $1701<m<$ 170 and $|m-l|\leq 40$. In the upper three panels of Fig. 13, we show the radial torque density for our three disc parameter combinations whilst adopting the particular value of $e=0.5$. This reference eccentricity is chosen since more noticeable discrepancies begin to arise between the linear framework and simulations beyond this value (see Fig. 6 for example). The convergence properties for this high value of $e$ will bound faster convergence behaviour for smaller values of $e$, whilst the larger values are expected to converge more slowly. In each panel, the coloured lines correspond to the profiles obtained with different maximal values of $|m-l|_{\rm max}$, whilst $m_{\rm max}=170$ is held fixed at the standard value. The results span from $|m-l|=5$ (dark purple line), up to the fiducial $|m-l|_{\rm max}=40$ (yellow line), in increments of 5. Note that we restrict our attention to the radial range $1.60.4<r/a<$ 1.6. Outside of this interval, the torque density is of lower amplitude and the torque wiggle features are predominantly formed by the action of low order modes in $m$ and $|m-l|$ (see discussion in section 4.2).
One clearly observes a change of the torque profiles as more off-diagonal $|m-l|$ modes are included. The maximum peak and minimum trough, located at peri and apocentre respectively, are underestimated and less well defined when fewer modes are included. Intuitively, Murray & Dermott (1999) show that the disturbing function Fourier components for eccentric perturbers are proportional to $e^{|m-l|}$. Therefore, for larger $e$, we expect the increasing importance of modal combinations with $l\neq m$. By the point that $|m-l|_{\rm max}\geq 35$ the profiles seem rather well converged, with the main features established and only very subtle changes as the included off-diagonal modes are expanded further.
@@
In Fig. 14 we now examine the convergence of the integrated torque for the full range of $e$. Akin to Fig. 6, we plot the net integrated torque curves as a function of $e$ for our three fiducial disc setups. Once again, the coloured lines denote the linear calculation results when including an increasing number of $|m-l|$ modes, from 5 (dark purple) up to 40 (yellow). All three panels tell a similar story. For sufficiently low values of $e$, well below the transonic transition and torque reversal, all of the curves overlap and a low number of off-diagonal modes satisfy convergence. Of course, for $e=0.0$ only the diagonal $m=l$ modes feature anyway. Pushing to higher values of $e$, we observe that the magnitude of the torque is severely underestimated unless a sufficient number of modes are included. The discrepancies are most pronounced just beyond the transonic turnover, where the net torque reaches an absolute minimum. It seems that for $|m-l|_{\rm max}=25$ and above, the curves all overlap and are sufficiently converged. Note that this convergence in the integrated net torque is slightly less stringent than the detailed convergence in the full radial torque density profile.
Once again, in order to further scrutinise the robustness of this fiducial modal range, we also plot the data points for the extended mode summation, as considered above for the disc with $p=0.5$ and $e=0.5$. The $|m-l|_{\rm max}=80$ and $m_{\rm max}=220$ markers closely overly the fiducial data point, once again indicating that the inclusion of a reasonable number of additional modes incurs no significant changes in the linear theory torque diagnostics. More quantitatively, radially integrating over the residuals $\Delta dT/dr$ seen in the lower panel of Fig. 13, gives a net discrepancy of $0.011$ and $-2.1\times 10^{-5}$ for the extended $|m-l|_{\rm max}$ and $m_{\rm max}$ runs respectively. Both of these corrections are still too small to bring the simulations and linear theory into perfect alignment which differ by $\sim 0.06$ for the fiducial modal sum at this specific data point.
Figure 14: Convergence study for the net integrated torque across our three disc profiles and $e$ values. The square markers and connective lines show the $T$ profiles as functions of $e$, for fixed $m_{\rm max}=170$ and increasing values of $|m-l|_{\rm max}$ from 5 (purple) to 40 (yellow). In the $p=0.5$ panel, we additionally plot the $e=0.5$ data point for the extended modal ranges with $(m_{\rm max},|l-m|_{\rm max})=(170,80)$ [red $\times$] and (220,40) [grey $+$].
@@ Appendix B A New Benchmark for planet-disc numerics
The focused study presented in this paper demonstrates the remarkable agreement that can be obtained between linear theory and numerical simulations, when appropriate care is taken. Having established this effective correspondence, we promote the use of our linear theory calculations as a controlled benchmark for testing future planet-disc interaction problems. Indeed, a detailed code versus theory comparison in the linear regime might also lend confidence to the numerical scheme as more general setups are considered (e.g. higher planet masses, different thermal prescriptions, viscous discs). To this end, we have provided a supplementary data repository which contains our torque density and AMF profiles across the three disc parameters combinations and a range of eccentricities, $e\in(0.01,0.06,0.12,0.3)$. This is accompanied by exemplar reading and plotting scripts which show how to access and visualise the data. Together, these offer a new standard benchmark for testing the performance of planet-disc problems. The repository can be accessed at https://github.com/callumf8/linear_ecc_planet.

### v2

@@ 5.2.2 Higher mass planets
Increasing the planetary mass further would result in more pronounced differences which we do not consider in this paper. Zhu & Zhang (2022) performed exemplary 2D simulations of eccentric Jupiter mass planets, showing that the spiral wake morphology is significantly modified compared with the low mass regime. Namely, the larger amplitude wakes can interact nonlinearly, leading to the messier interference of wake crossings. Furthermore, the rapid steepening and shocking of the waves deposits angular momentum more efficiently. This can promote gap opening and a transition towards Type II migration. Interestingly, as the eccentricity increases, the gap depth is found to be shallower (also see Duffell & Chiang, 2015) which might help constrain the properties of hypothesized planets in observed gaps.
@@ 5.2.4 Higher eccentricities
Whilst we push to highly supersonic values of $e$, we curtail our study at the upper value of $e=0.7$. By this point we are beginning to find small but noticeable discrepancies in the linear theory and simulation results. For example, the amplitude of sharp radial torque density features are not in exact agreement in Figs. 4-5 for $e=0.5$, whilst the linear theory seems to slightly underestimate the magnitude of numerical net torques in Fig. 7. However, in appendix A.1 we have demonstrated sufficient modal convergence in |ml||m-l| for the fiducial upper value of $m=170$ adopted by the linear framework. We also tested the use of expanded ranges with $m_{\rm max}=220$ and $|m-l|_{\rm max}=80$, but this did not provide a significant improvement on the comparison with simulations.
@@ Acknowledgements
CWF and AJD would like to thank the referee for a detailed report which helped improve the clarity of this work. The authors would also like to acknowledge insightful conversations with Roman Rafikov, Yoram Lithwick, and Eugene Chiang during the preparation of this manuscript. We also thank Pablo Benítez-Llambay for suggesting the inclusion of appendix B as a benchmark for future numerical investigations. This research was supported by the W. M. Keck Foundation Fund and the IAS Fund for Memberships in the Natural Sciences (CWF), and by NASA through the NASA Hubble Fellowship grant #HST-HF2-51553.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555 (AJD).
@@ A.1 Linear theory modal convergence
In this section, we will interrogate the modal convergence of our linear methodology. Indeed, since this paper considers highly eccentric orbits, beyond the scope which was confidently converged in the previous work of FR25, we anticipate the need for a high number of modes. Throughout the main body of this paper, we have assumed a default modal range of $1<m<170$ and $|m-l|\leq 40$. In the upper three panels of Fig. 14, we show the radial torque density for our three disc parameter combinations whilst adopting the particular value of $e=0.5$. This reference eccentricity is chosen since more noticeable discrepancies begin to arise between the linear framework and simulations beyond this value (see Fig. 7 for example). The convergence properties for this high value of $e$ will bound faster convergence behaviour for smaller values of $e$, whilst the larger values are expected to converge more slowly. In each panel, the coloured lines correspond to the profiles obtained with different maximal values of $|m-l|_{\rm max}$, whilst $m_{\rm max}=170$ is held fixed at the standard value. The results span from |ml|=5|m-l|=5 (dark purple line), up to the fiducial $|m-l|_{\rm max}=40$ (yellow line), in increments of 5. Note that we restrict our attention to the radial range $0.4<r/a<1.6$. Outside of this interval, the torque density is of lower amplitude and the torque wiggle features are predominantly formed by the action of low order modes in $m$ and |ml||m-l| (see discussion in section 4.2).
One clearly observes a change of the torque profiles as more off-diagonal |ml||m-l| modes are included. The maximum peak and minimum trough, located at peri and apocentre respectively, are underestimated and less well defined when fewer modes are included. Intuitively, Murray & Dermott (1999) show that the disturbing function Fourier components for eccentric perturbers are proportional to $e^{|m-l|}$. Therefore, for larger $e$, we expect the increasing importance of modal combinations with $l\neq m$. By the point that $|m-l|_{\rm max}\geq 35$ the profiles seem rather well converged, with the main features established and only very subtle changes as the included off-diagonal modes are expanded further.
@@
In Fig. 15 we now examine the convergence of the integrated torque for the full range of $e$. Akin to Fig. 7, we plot the net integrated torque curves as a function of $e$ for our three fiducial disc setups. Once again, the coloured lines denote the linear calculation results when including an increasing number of |ml||m-l| modes, from 5 (dark purple) up to 40 (yellow). All three panels tell a similar story. For sufficiently low values of $e$, well below the transonic transition and torque reversal, all of the curves overlap and a low number of off-diagonal modes satisfy convergence. Of course, for $e=0.0$ only the diagonal $m=l$ modes feature anyway. Pushing to higher values of $e$, we observe that the magnitude of the torque is severely underestimated unless a sufficient number of modes are included. The discrepancies are most pronounced just beyond the transonic turnover, where the net torque reaches an absolute minimum. It seems that for $|m-l|_{\rm max}=25$ and above, the curves all overlap and are sufficiently converged. Note that this convergence in the integrated net torque is slightly less stringent than the detailed convergence in the full radial torque density profile.
Once again, in order to further scrutinise the robustness of this fiducial modal range, we also plot the data points for the extended mode summation, as considered above for the disc with $p=0.5$ and $e=0.5$. The $|m-l|_{\rm max}=80$ and $m_{\rm max}=220$ markers closely overly the fiducial data point, once again indicating that the inclusion of a reasonable number of additional modes incurs no significant changes in the linear theory torque diagnostics. More quantitatively, radially integrating over the residuals $\Delta dT/dr$ seen in the lower panel of Fig. 14, gives a net discrepancy of $0.011$ and $2.1\times 10^{-5}$ for the extended $|m-l|_{\rm max}$ and $m_{\rm max}$ runs respectively. Both of these corrections are still too small to bring the simulations and linear theory into perfect alignment which differ by $\sim 0.06$ for the fiducial modal sum at this specific data point.
Figure 15: Convergence study for the net integrated torque across our three disc profiles and $e$ values. The square markers and connective lines show the $T$ profiles as functions of $e$, for fixed $m_{\rm max}=170$ and increasing values of $|m-l|_{\rm max}$ from 5 (purple) to 40 (yellow). In the $p=0.5$ panel, we additionally plot the $e=0.5$ data point for the extended modal ranges with $(m_{\rm max},|l-m|_{\rm max})=(170,80)$ [red $\times$] and (220,40) [grey ++].
@@ Appendix B A New Benchmark for planet-disc numerics
The focused study presented in this paper demonstrates the remarkable agreement that can be obtained between linear theory and numerical simulations, when appropriate care is taken. Having established this effective correspondence, we promote the use of our linear theory calculations as a controlled benchmark for testing future planet-disc interaction problems. Indeed, a detailed code versus theory comparison in the linear regime might also lend confidence to the numerical scheme as more general setups are considered (e.g. higher planet masses, different thermal prescriptions, viscous discs). To this end, we have provided a supplementary data repository which contains our torque density and AMF profiles across the three disc parameters combinations and a range of eccentricities, $e\in(0.01,0.06,0.12,0.3)$. This is accompanied by exemplary reading and plotting scripts which show how to access and visualise the data. Together, these offer a new standard benchmark for testing the performance of planet-disc problems. The repository can be accessed at https://github.com/callumf8/linear_ecc_planet.